# Research on Superalignment Should Advance Now with Parallel Optimization of Competence and Conformity

### v1

@@ 3 Paths towards Superalignment
Superalignment is achievable, with its core as constructing scalable high-quality supervision signals $\mathcal{S}\!=\!\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$ (Sec. 2.1). In terms of $\mathcal{S}$’s source, existing efforts fall into three paradigms: Sandwiching, Self-enhancement, and Weak-to-Strong Generalization, as shown in Fig. 2. We introduce each and analyze why they are insufficient to fully achieve superalignment, grounded in Sec. 2.
@@ 3.1 Sandwiching
Sandwiching (Cotra, 2021; Bowman et al., 2022) derives the solution $\mathbf{y}$ for each $\mathbf{x}$ that neither AI $\mathcal{A}$ nor a non-expert human $\mathcal{H}$ can solve alone through interactions. The underlying insight is that $\mathcal{A}$ has the necessary knowledge or skills that $\mathcal{H}$ lack, while $\mathcal{H}$ excels in understanding and judgment. Formally, $C(\mathcal{A})\!>\!\tau\!>\!C(\mathcal{H})$ and $\mathbb{E}_{\mathbf{x}}|U_{\mathcal{H}}(x)\!-\!U(x)|\!<\!\epsilon$, where $\tau$ is the minimal capacity for solving $\mathbf{x}$, $U_{\mathcal{H}}$ is the human’s own utility function (judgment) and $U$ is the golden one. In terms of the utilization of AI and the form of human signals, there are three main approaches:
(1) Consultancy (Khan et al., 2024; Kenton et al., 2024): $\mathcal{A}$ provides relevant information, reasoning, or candidate solution $\mathbf{y}$, and then $\mathcal{H}$ judges $\mathbf{y}$ by $U_{\mathcal{H}}(\mathbf{y})$, offers feedback, and asks $\mathcal{A}$ to refine $\mathbf{y}$ accordingly until convergence.
(2) Debate (Irving et al., 2018; Brown-Cohen et al., 2023; Du et al., 2023b; Kenton et al., 2024; Kirchner et al., 2024; Khan et al., 2024): This method involves multiple AI systems, $\{\mathcal{A}_{1},\dots,\mathcal{A}_{K}\}$. Each $\mathcal{A}_{i}$ of them provides a (probably different) candidate solution $\mathbf{y}_{i},i\!=\!1,\dots,K$, and tries to defend/justify $\mathbf{y}_{i}$ while criticizing others. The human $\mathcal{H}$ then interacts with them and determines the most appropriate solution. Compared to Consultancy, Debate introduces two relaxations: i) Judging the defense: $\mathcal{H}$ assesses $U(\mathbf{y}_{i})$ indirectly by evaluating AI debaters’ rationale. ii) Relative comparison: Instead of estimating absolute $U(\mathbf{y}_{i})$, $\mathcal{H}$ conducts relative comparison, which only requires $\left|[U_{\mathcal{H}}(\mathbf{y}_{i})\!-\!U_{\mathcal{H}}(\mathbf{y}_{j})]\!-\![U(\mathbf{y}_{i})\!-\!U(\mathbf{y}_{j})]\right|\!<\!\epsilon$ and thus is more reliable (Mohankumar & Khapra, 2022).
@@
(3) Recursive Amplification (Christiano et al., 2018; Charbel-Raphaël, 2023; Radhakrishnan et al., 2023)  This method requires $\mathcal{H}$ to decompose a complex task $\mathbf{x}$ into subtasks $\mathbf{x}\!=\!(\mathbf{x}^{1},\dots,\mathbf{x}^{K})$. Then each $\mathcal{A}_{i}$ (or the same one) provides a partial solution $\mathbf{y}^{i}$ for $\mathbf{x}^{i}$ and the human $\mathcal{H}$ aggregates them to derive the final solution $\mathbf{y}$. The relaxation lies in that human can judge the partial solution better than the complete one, i.e., $|U_{\mathcal{H}}(\mathbf{y}^{i})\!-\!U(\mathbf{y}^{i})|\!<\!|U_{\mathcal{H}}(\mathbf{y})\!-\!U(\mathbf{y})]|,\forall i$. Such a schema can be repeated across multiple rounds with $\mathcal{A}_{i}$ improved by the derived $(\mathbf{x},\mathbf{y})$ in each round, and applied to specific downstream tasks (Sturgeon et al., 2023; Wen et al., 2024) and reward modeling (Leike et al., 2018).
@@
Table 1: Debate performance evaluated on QuALITY (Pang et al., 2022) and GPQA (Rein et al., 2024). Helpful Debater: $\mathcal{A}$ stands for the ground truth answer. Sneaky Debater: $\mathcal{A}$ defends an alternative (wrong) answer. We find (1) a stronger sneaky debater can mislead the judge ((a) v.s. (b)) and (2) a stronger judge is more important for accuracy than a stronger debater ((f) v.s. (c)). Full results and additional analysis are in Tab. A.4, Sec. A.2.2 and  A.3.1.
@@
Limitations:  Despite promising results, this paradigm has fundamental limitations. (1) Reliance on Human Capability: Humans must be able to objectively evaluate AI-generated arguments, insights, or partial solutions and effectively handle task decomposition and integration, when they cannot directly assess or provide the final solution. See empirical evidence in Tab. 3.1 which shows that humans might be misled by AI debaters. (2) Task Requirements: The task must be understandable to humans (for identifying knowledge or task decomposition), decomposable (for recursive amplification), and at least partially solvable by AI (e.g., completing sub-tasks). (3) Supervision Signal Noise: Uncertainty in both humans and AI (Testoni & Fernández, 2024) means optimal solutions cannot always be elicited. Besides, biases in both humans and AI, along with AI’s sycophancy (Sharma et al., 2024b), can lead to over-alignment with unreliable human feedback or deception (Hagendorff, 2024)—especially when humans fail to detect AI’s flaws.
Argument: We claim that Sandwiching may not fully achieve superalignment due to the three fundamental limitations described above. Especially for ASI and the task $\mathbf{x}$ beyond human’s scope, that is, $C(\mathcal{A}_{\text{ASI}})\!>\!\tau\!\gg\!C(\mathcal{H})$, humans are likely to fail in task decomposition or utility judgment, i.e., $|U(\mathcal{H}(x))\!-\!U(\mathcal{A}_{\text{ASI}}(x))|$ as described in Premise 2 in Sec. 2.2 and $|U_{\mathcal{H}}(x)\!-\!U(x)|$ could both be large.
@@ 3.2 Self-Enhancement
Self-enhancement (Bai et al., 2022c; Huang et al., 2024) enables AI $\mathcal{A}$ to improve self-generated solution $\mathbf{y}$ based on feedback from itself or multiple AI with similar capability (by self-revising or searching), involving minimal or no external human supervision. This paradigm can be traced back to the self-play paradigm (Silver et al., 2017), a.k.a., Expert Iteration (Anthony et al., 2017) originally adopted in the famous AlphaZero system. Based on AI involved and enhancement methods, there are two approaches:
@@
(1) Self-Refinement (Saunders et al., 2022; Ganguli et al., 2023): A single system, $\mathcal{A}$, performs an iterative generate-then-refine process. At iteration $t$, it produces a candidate $\mathbf{y}^{t}$, then progressively self-critiques and refines it into $\mathbf{y}^{t+1}$, and repeats until convergence. This can occur without model training, known as In-Context Alignment (ICA) (Lin et al., 2023; Gou et al., 2024), or through alternating refinement and fine-tuning, where $\mathbf{y}^{t+1}$ updates $\mathcal{A}^{t}$ to form $\mathcal{A}^{t+1}$ for the next iteration — a process known as Learning from AI Feedback (LAIF) (Gulcehre et al., 2023; Lee et al., 2024).
(2) Collective Enhancement: Multiple $\mathcal{A}_{i}$ interact with each other to arrive at a final $\mathbf{y}$. The underlying idea is that if multiple approaches converge on the same solution, confidence in it increases (Wang et al., 2023a). This can be done by Debate without human judge (multiple debaters reach a consensus) (Chern et al., 2024), majority voting among independent $\mathcal{A}_{i}$ (Luo et al., 2023), or social simulations where each $\mathcal{A}_{i}$ takes a particular role (Pang et al., 2024).
@@
Limitations:  Similarly, several limitations also lie in the Self-Enhancement paradigm. (1) Error Accumulation (Du et al., 2023a): Since supervision signals are entirely derived by AI, any errors or noise can propagate and accumulate over time, potentially leading to issues like polarization (Piao et al., 2025; Lai et al., 2024). Tab. A.4 in Appendix shows that continual training on self-generated data progressively degrades performance as model size increases. (2) Self-Preference (Wataoka et al., 2024): Because AI is inherently biased towards self-generated content, it results in suboptimal critique and verification performance (false positives). (3) Mode Collapse (Shumailov et al., 2024): Iterative fitting on self-generated data leads to degeneration into a restricted high-frequency sub-distribution, resulting in a loss of generative diversity and associated capability.
Argument: We argue that this paradigm’s effectiveness is quite limited, for two reasons: i) Insufficient Capacity. Self-enhancement doesn’t increase capacity $C(\mathcal{A})$, negating Premise 1 in Sec. 2.2. This can be regarded as a form of self-training (Feng et al., 2023) minimizing $\text{KL}[\mathcal{A}^{t-1}(\mathbf{x},\mathbf{y})||\mathcal{A}^{t}(\mathbf{x},\mathbf{y})]$, where KL is the Kullback–Leibler divergence and $\mathcal{A}^{t}(\mathbf{x},\mathbf{y})$ is the internal data distribution in $\mathcal{A}$ at iteration $t$, which elicits and reinforces latent capability but cannot expand it, as it is upper bounded by the capacity (Sharma et al., 2024a). ii) Mode Collapse might even decrease model capacity (Seddik et al., 2024).
@@ 3.3 Weak-to-Strong Generalization
Weak-to-Strong Generalization (W2SG) (Burns et al., 2023): This paradigm involves $K$ models, $\mathcal{A}_{1},\!\dots\!,\mathcal{A}_{K}$, with increasing capacities $C(\mathcal{A}_{1})\!<\!\dots\!<\!C(\mathcal{A}_{K})$, where $C(\mathcal{A}_{K})\!=\!C(\mathcal{A}_{\text{ASI}})$, to solve task $\mathbf{x}$ that requires superintelligence. The goal is to have each $\mathcal{A}_{i}$ generate a solution $\mathbf{y}_{i}$, then fine-tune $\mathcal{A}_{i+1}$ using $\mathbf{y}_{i}$, $i\!=\!1,\!\dots\!,K\!-\!1$ in a progressive manner (called bootstrapping). Therefore, each $\mathcal{A}_{i+1}$ surpasses its predecessor $\mathcal{A}_{i}$. The key insight is to bridge the internal knowledge $C(\mathcal{A}_{i})$ gained through pretraining with the demanded capability of the target task $U(\mathbf{x})$, eliciting latent knowledge (Mallen et al., 2024) from more advanced models rather than teaching them entirely new skills. Below are three key directions:
(1) Vanilla W2SG (Burns et al., 2023; Sang et al., 2024; Liu & Alahi, 2024; Yang et al., 2024c): The weaker $\mathcal{A}_{i-1}$ with lower capacity generates a solution $\mathbf{y}_{i-1}$ to supervise a stronger $\mathcal{A}_{i}$. Performance can be improved by adding an auxiliary confidence loss or incorporating model ensemble.
(2) Co-Generalization (Lyu et al., 2024): Both the weaker $\mathcal{A}_{i-1}$ and the stronger $\mathcal{A}_{i}$ jointly produce synthetic solutions $\mathbf{y}$, so that $\mathcal{A}_{i}$’s capacity is fully utilized in providing better and less noisy supervision signals.
(3) Reward Model Generalization (Sun et al., 2024): Not the policy model $\mathcal{A}$, but the reward model is supervised in the W2SG schema, grounded in the intuition that evaluation is easier than generation (Naor, 1996). The supervised reward model then provides signals for the policy model with the same capacity in the subsequent alignment process.
@@
Figure 3: Training accuracy of Qwen2.5-7B-Instruct on CosmosQA with different ratios (0.0 to 1.0) of noise (flipped labels) in the training set. The model shows a stronger tendency to fit noise due to its large parameter size, except in the range of 0.4 to 0.6 (Green), where labels are heavily randomized. In contrast, Fig. 5 shows that smaller models struggle more with noise. More results and discussions are presented in Tab. A.4 and Tab. LABEL:tab:noise_sciq in Appendix.
@@
Limitations: While W2SG is relatively practical and requires no human intervention, relying on synthetic signals from weaker $\mathcal{A}_{i-1}$ to supervise stronger $\mathcal{A}_{i}$ can lead to unintended behavior, particularly when there is a significant gap between $\mathcal{A}_{i-1}$ and $\mathcal{A}_{i}$. This stems from two problems: (1) Imitation of weak supervision (Burns et al., 2023). The stronger $\mathcal{A}_{i+1}$ may merely imitate how the weaker $\mathcal{A}_{i}$ behave, resulting in a perfect replication of $\mathcal{A}_{i}$ with limited improvement gap (Christiano et al., 2022) instead of generalization. Besides, strong $\mathcal{A}_{i+1}$ tends to overfit errors/bias from $\mathcal{A}_{i}$ when synthetic signals contain systematic errors, hurting performance. As shown in Fig. 3, we observe that a highly capable AI is more susceptible to such errors. (2) Task Salience in the Strong AI: W2SG is effective only when the task $\mathbf{x}$ being elicited is internally “salient” to $\mathcal{A}_{i+1}$. The evidence lies in that in tasks where W2SG succeeds, simple 5-shot prompting also performs comparably (Burns et al., 2023). Eliciting non-salient capabilities is more challenging. For safety, it’s important that risks caused by current AI may largely differ from those exhibited by future ASI.
Argument:  We believe that the current W2SG practice is insufficient for achieving superalignment. From the limitations discussed above, the key issue is the gap between the weak $\mathcal{A}_{i-1}$ and strong $\mathcal{A}_{i}$, $\delta=|U(\mathcal{A}_{i})-U(\mathcal{A}_{i-1})|$. Imitation of weak supervision occurs when $\delta$ is large. Although (Christiano et al., 2022) shows that larger $\mathcal{A}_{i}$ agrees less with the errors of smaller $\mathcal{A}_{i-1}$ (or this can be reinforced through confidence loss), in superintelligence-level tasks $\mathbf{x}$, it is infeasible to determine whether $\mathcal{A}_{i}$ or $\mathcal{A}_{i-1}$ is correct ($\mathcal{A}_{i}$’s self-confidence may be biased). Tab. A.4 in Appendix demonstrates this: when all synthetic labels are flipped, $\mathcal{A}_{i}$ achieves 100% training accuracy but poor test accuracy. When the gap $\delta$ is too large, it violates Premise 2 in Sec. 2.2, causing discussed issues, while a small $\delta$ leads to the Task Salience problem, where the difficult task is not salient for $\mathcal{A}_{i}$, hindering its ability to uncover novel solutions.
4 A Novel Conceptual Framework

### v2

@@ 3 Existing Paths towards Superalignment
To support arg.1, i.e., superalignment is achievable, we begin by examining why it is viewed as infeasible. Since superalignment is formulated as eliciting latent capabilities Christiano et al. (2021) and closing the gap between $U(\mathcal{A})$ and $C(\mathcal{A})$, we present two premise necessary for the feasibility of superalignment:
  • Premise 1: Superhuman capacity, $C(\mathcal{A})\!\gg\!C(\mathcal{H})$, can be achieved through model scaling (larger models and more training data) to gain information (Brown et al., 2020), knowledge (Ovadia et al., 2024), and skills (Sharma et al., 2022).
  • Premise 2: If $C(\mathcal{A})$$U(\mathcal{A})$ are sufficiently high (Wei et al., 2022a), $C(\mathcal{A})$ $C(\mathcal{H})$, $U(\mathcal{A})\!<\!U(\mathcal{H})$, then moderate gaps, $\delta_{\mathcal{A},\mathcal{H}}\!=\!|\mathbb{D}_{U}[U(\mathcal{A}),U(\mathcal{H})]\!-\!\mathbb{D}_{C}[C(\mathcal{A}),C(\mathcal{H})]|$, can be closed without golden signals.
We argue that the perceived infeasibility lies in that current paradigms fail to meet these conditions. Therefore, we analyze their fundamental limitations to motivate solutions and support superalignment feasibility. Grounded in Sec. 2.1, the core of achieving Def. 2.1 is to construct scalable high-quality supervision signals $\mathcal{S}\!=\!\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$. In terms of $\mathcal{S}$’s source, existing efforts fall into three paradigms: Sandwiching, Self-enhancement, and Weak-to-Strong Generalization, as shown in Fig. 2.
@@ 3.1 Sandwiching
For a $\mathbf{x}$ that neither AI $\mathcal{A}$ nor non-expert human $\mathcal{H}$ can solve alone, Sandwiching (Cotra, 2021; Bowman et al., 2022) derives $\mathbf{y}$ through interactions. The insight is that $\mathcal{A}$ has the necessary knowledge or skills that $\mathcal{H}$ lack, while $\mathcal{H}$ excels in understanding and judgment. Formally, $C(\mathcal{A})\!>\!\tau_{\mathbf{x}}\!>\!C(\mathcal{H})$ and $\mathbb{E}_{\mathbf{x}}|\hat{U}_{\mathcal{H}}(\mathbf{x})\!-\!U(\mathbf{x})|\!<\!\epsilon$, where $\tau_{\mathbf{x}}$ is the minimal capacity for solving $\mathbf{x}$, $\hat{U}_{\mathcal{H}}$ is the human’s own utility function (judgment, approximating $U$). In terms of how to use AI and form human signals, there are three main approaches:
@@
(1) Consultancy Khan et al. (2024); Kenton et al. (2024): $\mathcal{A}$ provides relevant information, reasoning, or candidate solution $\mathbf{y}$, and then $\mathcal{H}$ judges $\mathbf{y}$ by $\hat{U}_{\mathcal{H}}(\mathbf{y})$, offers feedback, and asks $\mathcal{A}$ to refine $\mathbf{y}$ accordingly until convergence. (2) Debate Irving et al. (2018); Brown-Cohen et al. (2023); Du et al. (2023b); Kenton et al. (2024); Kirchner et al. (2024); Khan et al. (2024): Each $\mathcal{A}_{k}$ from a set of AI systems, $\{\mathcal{A}_{k}\}_{k=1}^{K}$, provides a different candidate solution $\mathbf{y}_{k}$, and tries to defend $\mathbf{y}_{i}$ while criticizing others. Then human $\mathcal{H}$ interacts with them to determine the best one. Debate introduces two relaxations to Consultancy: i) $\mathcal{H}$ estimates $\mathbf{y}_{i}$’s golden utility indirectly by examining AI debaters’ rationale with $\hat{U}_{\mathcal{H}}$. ii) $\mathcal{H}$ conducts relative comparison, which only requires $\left|[\hat{U}_{\mathcal{H}}(\mathbf{y}_{i})\!-\!\hat{U}_{\mathcal{H}}(\mathbf{y}_{j})]\!-\![U(\mathbf{y}_{i})\!-\!U(\mathbf{y}_{j})]\right|\!<\!\epsilon$ and hence is much more reliable (Mohankumar and Khapra, 2022). (3) Recursive Amplification Christiano et al. (2018); Charbel-Raphaël (2023); Radhakrishnan et al. (2023)  $\mathcal{H}$ decomposes a complex task $\mathbf{x}$ into subtasks $\{\mathbf{x}^{k}\}_{k=1}^{K}$. Then each $\mathcal{A}_{i}$ provides a partial solution $\mathbf{y}^{i}$ for $\mathbf{x}^{i}$ and $\mathcal{H}$ aggregates them to derive the final solution $\mathbf{y}$. The relaxation lies in that humans can judge the partial solution better than the complete one. Such a schema can be repeated across multiple rounds with $\mathcal{A}_{i}$ improved by the derived $(\mathbf{x},\mathbf{y})$ in each round, and applied to specific downstream tasks (Sturgeon et al., 2023; Wen et al., 2024) and reward modeling (Leike et al., 2018).
@@
Table 1: Debate performance on QuALITY (Pang et al., 2022) and GPQA (Rein et al., 2024). Helpful Debater stands for the ground truth answer. Sneaky Debater defends an alternative (wrong) answer. We find (1) a stronger sneaky debater can mislead the judge ((a) v.s. (b)) and (2) a stronger judge is more important than a stronger debater ((f) v.s. (c)). Full results and analysis are shown in Tab. A.9, Sec. A.7.2 and  A.8.1.
@@
Limitations:  Despite promising results, Sandwiching has fundamental limitations. (1) Reliance on Human Capability: Humans must be able to objectively evaluate AI-generated arguments, insights, or partial solutions and effectively handle task decomposition and integration. However, for superhuman tasks $\mathbf{x}$, the gap $|\hat{U}_{\mathcal{H}}(\mathbf{y})-U(\mathbf{y})|$ could be quite large, especially when we consider the high uncertainty and noise inherent in both humans and AI (Testoni and Fernández, 2024), e.g., AI’s sycophancy (Sharma et al., 2024b) and deception (Hagendorff, 2024). This leads to unreliable supervision $\mathcal{S}$ and makes Sandwiching fail to meet Premise 2. See empirical evidence in Tab. 3.1, which shows that humans might be misled by AI debaters. (2) Task Requirements: The task $\mathbf{x}$ must be understandable to humans (for judgment or decomposition), decomposable (for recursive amplification), and at least partially solvable by AI. However, this does not hold as superhuman tasks $\mathbf{x}$ are naturally beyond human scope, i.e., $C(\mathcal{A}_{\text{ASI}})\!>\!C(\mathcal{A})\!\gg\!C(\mathcal{H})$. Without altering AI models, Sandwiching fails Premise 1.
@@ 3.2 Self-Enhancement
Self-enhancement (Bai et al., 2022c; Huang et al., 2024) enables AI $\mathcal{A}$ to improve self-generated solution $\mathbf{y}$ based on feedback from itself or multiple AI with similar capability, involving minimal external human supervision. This paradigm can be traced back to self-play (Silver et al., 2017), a.k.a., Expert Iteration (Anthony et al., 2017) originally adopted in AlphaZero. Based on AI involvement and enhancement methods, we discuss two major approaches:
@@
(1) Self-Refinement (Saunders et al., 2022; Ganguli et al., 2023): A single system, $\mathcal{A}$, performs an iterative generate-then-refine process. At iteration $t$, it produces a candidate $\mathbf{y}^{t}$, then self-critiques and refines it into $\mathbf{y}^{t+1}$, repeating until convergence. This can occur without model training, known as In-Context Alignment (ICA) (Lin et al., 2023; Gou et al., 2024), or through alternating refinement and fine-tuning, where $\mathbf{y}^{t+1}$ updates $\mathcal{A}^{t}$ to form $\mathcal{A}^{t+1}$ for the next iteration — a process known as Learning from AI Feedback (LAIF) (Gulcehre et al., 2023; Lee et al., 2024). (2) Collective Enhancement: Multiple $\mathcal{A}_{i}$ interact with each other and arrive at a final $\mathbf{y}$. If multiple approaches converge on the same solution, confidence in it would increase (Wang et al., 2023a). This can be done by Debate without a human judge (multiple debaters reach a consensus) (Chern et al., 2024), majority voting among independent $\mathcal{A}_{i}$ (Luo et al., 2023), or social simulations where each $\mathcal{A}_{i}$ takes a particular role (Pang et al., 2024).
@@
Limitations:  Similarly, several limitations also lie in the Self-Enhancement paradigm. i) Insufficient Capacity. This paradigm can be regarded as a form of self-training Feng et al. (2023) minimizing $\text{KL}[\mathcal{A}^{t-1}(\mathbf{x},\mathbf{y})||\mathcal{A}^{t}(\mathbf{x},\mathbf{y})]$, where KL is the Kullback–Leibler divergence and $\mathcal{A}^{t}(\mathbf{x},\mathbf{y})$ is the internal data distribution in $\mathcal{A}$ at iteration $t$, which only elicits and reinforces $C(\mathcal{A})$ but cannot expand it (Sharma et al., 2024a). Worse still, this process could cause mode collapse Shumailov et al. (2024): iterative fitting on self-generated data leads to degeneration into a restricted high-frequency sub-distribution, even losing capacity and negating Premise 1. ii) Damaged Capability. In such an iterative refinement process, where $\mathcal{A}_{i-1}$ evaluates the synthetic data for its successor $\mathcal{A}_{i}$, the estimated utility would be biased, i.e., $\hat{U}_{\mathcal{A}_{i-1}}(\mathcal{A}_{i}(\mathbf{y}))\!>\!U(\mathcal{A}_{i}(\mathbf{y}))$, known as self-preference Wataoka et al. (2024). Besides, since supervision signals $\mathcal{S}$ are entirely derived by AI, any errors or noise can propagate and accumulate over time, leading to issues like error accumulation (Du et al., 2023a) and polarization (Piao et al., 2025; Lai et al., 2024), and losing generative diversity as well as associated capability. Tab. LABEL:tab:aif_result in Appendix shows that continual training on self-generated data progressively degrades performance as model size increases. As a result, the system optimizes for self-preference rather than objective utility, and thus, even if $C(\mathcal{A})$ remains, the capability-capacity gap $|\mathbb{D}_{U}-\mathbb{D}_{C}|$ cannot be reduced, or would even increase, undermining Premise 2.
@@ 3.3 Weak-to-Strong Generalization
Figure 3: Training accuracy of Qwen2.5-7B-Instruct on CosmosQA with different ratios of noisy (flipped labels) training samples. LLMs show a stronger tendency to fit noise, except in the range of 0.4 to 0.6 (Green), where labels are heavily randomized. In contrast, Fig. 5 shows that smaller models struggle more with noise. More results are in Tab. A.9 and Tab. LABEL:tab:noise_sciq.
Weak-to-Strong Generalization (W2SG) (Burns et al., 2023): This paradigm also involves $K$ models but with increasing capacities $C(\mathcal{A}_{1})\!<\!\dots\!<\!C(\mathcal{A}_{K})$, where $C(\mathcal{A}_{K})\!=\!C(\mathcal{A}_{\text{ASI}})$, to solve task $\mathbf{x}$ that requires superintelligence. The goal is to have each $\mathcal{A}_{k}$ generate a solution $\mathbf{y}_{k}$, then fine-tune $\mathcal{A}_{k+1}$ using $\mathbf{y}_{k}$, $k\!=\!1,\!\dots\!,K\!-\!1$ in a progressive manner (called bootstrapping). Therefore, each $\mathcal{A}_{k+1}$ surpasses its predecessor $\mathcal{A}_{k}$. The key is to bridge internal knowledge $C(\mathcal{A}_{k})$ with the capability demanded of the target task, eliciting latent knowledge Mallen et al. (2024) from more advanced models rather than teaching them entirely new skills. Below are three key directions: (1) Vanilla W2SG Burns et al. (2023); Sang et al. (2024); Liu and Alahi (2024); Yang et al. (2024c): The weaker $\mathcal{A}_{i-1}$ with lower capacity generates $\mathbf{y}_{i-1}$ to supervise a stronger $\mathcal{A}_{i}$. Performance can be improved with an auxiliary confidence loss or a model ensemble. (2) Co-Generalization (Lyu et al., 2024): Both the weaker $\mathcal{A}_{i-1}$ and the stronger $\mathcal{A}_{i}$ jointly produce synthetic solutions $\mathbf{y}$, so that $\mathcal{A}_{i}$’s capacity is fully utilized in providing better and less noisy supervision signals. (3) Reward Model Generalization (Sun et al., 2024): Not the policy model $\mathcal{A}$, but the reward model is supervised in the W2SG schema, grounded in the intuition that evaluation is easier than generation (Naor, 1996). The supervised reward model then provides signals for the policy model in the subsequent alignment process.
@@
Limitations: While W2SG is relatively practical, relying on synthetic signals from weaker $\mathcal{A}_{i-1}$ can lead to unintended behavior, particularly when there is a significant gap $\Delta^{(U)}\!=\!|U(\mathcal{A}_{i})\!-\!U(\mathcal{A}_{i-1})|$. This stems from two problems: (1) Imitation of Weak Supervision (Burns et al., 2023). The stronger $\mathcal{A}_{i+1}$ merely imitates or overfits the weaker $\mathcal{A}_{i}$’s behaviors and errors, resulting in a perfect replication of $\mathcal{A}_{i}$ with limited improvement gap Christiano et al. (2022) instead of generalization. As shown in Fig. 3, we observe that a highly capable AI is more susceptible to such errors. Although Christiano et al. (2022) shows that larger $\mathcal{A}_{i}$ agrees less with smaller $\mathcal{A}_{i-1}$’s errors, in superintelligence-level tasks, it’s infeasible to determine whether $\mathcal{A}_{i}$ or $\mathcal{A}_{i-1}$ is correct. Tab. A.9 demonstrates this: when all synthetic labels are flipped, $\mathcal{A}_{i}$ achieves 100% training accuracy but poor test accuracy. This imitation occurs when $\Delta^{(U)}$ is large, invalidating Premise 2. (2) Task Salience in the Strong AI: W2SG is effective only when the task $\mathbf{x}$ being elicited is internally “salient" to $\mathcal{A}_{i+1}$. The evidence lies in that in tasks where W2SG succeeds, simple 5-shot prompting also performs comparably Burns et al. (2023). Eliciting non-salient capabilities is more challenging. A small $\delta$ leads to this problem, where the difficult task is not salient for $\mathcal{A}_{i}$, violating Premise 1.
We can find that existing methods fail to meet both Premises 1 & 2. We argue that solutions exist and detail them in Sec. 4. More analyses of these limitations are provided in Appendix A.6.
4 Alternating Optimization Framework
In Sec. 3, we analyzed that existing paradigms fail to meet Premise 1 & Premise 2 by identifying their limitations. Moreover, they focus solely on task competence while overlooking unpredictable risks, especially when sensitive scenarios $\mathbf{x}$ are unknown. To address this, we first argue that, beyond existing paradigms, solutions enabling superalignment’s feasibility do exist (Sec. 4.1), introduce two essential principles critical to satisfying the premises (Sec. 4.2), and then present ACCO, a conceptual framework that operationalizes these principles (Sec. 4.3), jointly optimizing competence and conformity (arg.3), as illustrated in Fig. 4, and explore potential methods of evaluation (Sec. 4.4).